OpenStudy (anonymous):

Find the partial derivatives of n=0.04(T/P) when P=40 and T=600

OpenStudy (anonymous):

I know that when finding the partial with respect to P, T is treated as a constant. When I try it, I get n=0.04(600/1) =24

OpenStudy (anonymous):

Excuse me, \(n=0.04\frac{T}{P}\)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yeah, I think you are right

OpenStudy (anonymous):

I know the answer is -0.015 I just don't know how to get there.